OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6

OpenStudy (zpupster):

here maybe use substitution from first equation x+3y = -1 subtract 3y from both sides x-3y-3y = -1-3y x=-1-3y we know x so substitute in 2nd equation to find y 2(-1-3y) + 2y = 6 multiply thru parentheses -2-6y + 2y = 6 combine the like terms -2-4y=6 add 2 both sides -2+2-4y = 6+2 -4y=8 can you get it from here?